### v3

# LieRE: Lie Rotational Positional Encodings

Abstract
Transformer architectures rely on position encodings to capture token dependencies. Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations. However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity. We address these challenges with Lie Relative Encodings (LieRE), which replaces RoPE’s block-2D rotation matrix with a learned, dense, high-dimensional rotation matrix of variable density. Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 2% relative improvement over state-of-the-art baselines on 2D tasks and 1.5% on 3D tasks, while demonstrating superior generalization to higher resolutions. Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100, and we release our code to facilitate further research.
Machine Learning, ICML
1 Introduction
The attention mechanism, particularly within transformer architectures, has revolutionized machine learning across diverse domains. However, attention is inherently permutation—invariant—it cannot leverage the sequential order of its inputs as information. This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language processing to image understanding (Vaswani et al., 2017).
This challenge has sparked extensive research into positional encodings, which inject order information into the otherwise order-agnostic attention mechanism. The field has evolved from early approaches using fixed sinusoidal positional embeddings (Vaswani et al., 2017) to more sophisticated learned embeddings (Shaw et al., 2018a; Devlin et al., 2019; Dosovitskiy et al., 2020). Recent advances have introduced increasingly dynamic methods, including relative position representations (Shaw et al., 2018b; Dosovitskiy et al., 2020) and rotary positional embeddings (Su et al., 2024; Heo et al., 2024), demonstrating the critical role of position encoding in enabling transformers to effectively process ordered sequences.
In particular, Rotary Position Encoding (RoPE) has emerged as an elegant solution for encoding relative positional information in transformers (Su et al., 2024). RoPE works by applying a rotation matrix to each token’s keys and queries, where the rotation angle depends on the token’s absolute position in the sequence. The key insight is that when two tokens interact through the attention inner product, their rotated representations naturally encode their relative distance. For example, when a token at position five attends to a token at position two, RoPE enables the model to understand that these tokens are three positions apart, regardless of their absolute positions in the sequence. This translation-invariant property makes RoPE particularly efficient at capturing position-dependent patterns in text, which has led to its adoption in popular open-source language models such as LLaMA and Mixtral.
Figure 1: Overview of LieRE, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$. $d$ is the head dimension. The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
Despite RoPE’s success in sequential tasks (Touvron et al., 2023; Chowdhery et al., 2023), it faces several fundamental limitations. First, RoPE is designed for one-dimensional sequence data with an exponential decay in frequencies, making it suboptimal for spatial relationships in images or spatiotemporal patterns in videos where distant pixels may have strong correlations. Second, when handling higher dimensional data, the challenge compounds significantly–learned relative position encodings must capture exponentially many relative positions for $n$-dimensional data, as demonstrated in 2D image data where positions must be encoded both horizontally and vertically (Shaw et al., 2018a). Third, RoPE’s reliance on sparse, block-diagonal rotation matrices with handcrafted basis functions constrains its ability to learn complex spatial dependencies (Chu et al., 2024).
Our key insight is that Lie group theory provides a natural framework for generalizing relative position encodings to higher dimensions. We introduce Lie Relative Encodings (LieRE), which replaces RoPE’s rotation matrices with learned, dense rotations derived from Lie group generators. LieRE learns a basis of skew-symmetric matrices ${A_{i}}$, computes rotation matrices $R(p)=\exp(\sum_{i=0}^{n}p_{i}A_{i})$ for $n$-dimensional positions, and applies these rotations to keys and queries in the attention mechanism. The relative positions are then naturally captured through the inner product of the rotated keys and queries.
This approach addresses RoPE’s limitations in two key ways: (1) it handles higher-dimensional spatial relationships through Lie groups and (2) it increases representational capacity through variably dense, learned rotation matrices while requiring only 0.68% of parameters in a ViT-B model. Importantly, LieRE can be implemented with a single modification to existing architectures.
To assess the impact of LieRE and other positional encodings on ViT performance, we evaluate several encoding schemes across diverse tasks, including 2D and 3D image classification. Additionally, we investigate a fundamental spatial reasoning task where models must identify arrow directions. Despite their sophistication, contemporary multimodal LLMs struggle with this seemingly simple task. Our experiments reveal that successful completion of this task specifically requires relative position encodings, highlighting their crucial role in spatial understanding.
Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed (Heo et al., 2024), VisionLlama (Chu et al., 2024). First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix. Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful). MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
2 Related Work
2.1 Position Encodings
We organize our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a tex or top left corner of an image. Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the input. Methods such as sinusoidal and learned absolute encodings directly add vectors to the input token embedding (Vaswani et al., 2017; Devlin et al., 2019; Dosovitskiy et al., 2020).
Relative position encodings in contrast encode the relative positions of two tokens. One strategy is to learn an set of embeddings for position deltas which can be incorporated into the attention mechanism (Shaw et al., 2018b; Liu et al., 2021, 2022). However, this incurs quadratic computational cost in terms of the number of tokens as a separate embedding is required for every pair of tokens. Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product. The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism (Su et al., 2024). RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models (Touvron et al., 2023; Chowdhery et al., 2023; Jiang et al., 2024). However, RoPE can perform poorly on inference for larger context sizes than the model was trained on. This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
We refer to the last category of positional encodings as contextual position embeddings. This category is defined by encodings that aim to capture semantic positional information lost in traditional absolute and relative position encodings, often motivated by reasoning or mathematical tasks. Contextual Position encodings achieve (CoPE) this by allowing the model to learn how the position is computed (Golovneva et al., 2024). Abacus embeddings enable transformers to learn how to handle arithmetic by better exposing the digit structure of numbers (McLeish et al., 2024).
2.2 Extensions of RoPE
The efficiency and popularity of RoPE have led to several lines of work building off of it.
One notable one is context extension, which aims to address the fact that RoPE NLP models trained on short documents tend to perform poorly on long documents. Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning (Ding et al., 2024; Peng et al., 2023; Tworkowski et al., 2024; Chen et al., 2023).
The final, and most relevant, line of work has been specifically focused on adapting RoPE to image tasks. Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE capable of encoding 2D positional data (Chu et al., 2024; Heo et al., 2024). The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
3 Background
3.1 Lie Groups in the Context of Attention
In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE. Lie groups are well studied, especially in the context of representation theory, and standard texts including (Fulton & Harris, 2013) are able to provide a more extensive introduction to the subject.
In this context, Lie groups are smooth sets of matrices that are closed under matrix multiplication and inversion. For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$, also known as the Lie Algebra, to the Lie group. The exponential map is a diffeomorphism and has the following key property for $U,V\in\mathbb{R}^{n\times n}$ close together:
$\exp(U-V)=\exp(-V+U)\approx\exp(V)^{-1}\exp(U)$ (1)
Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 4. These form a special Lie group that is commutative, allowing us to strengthen the statement in (1) to
$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(V)^{-1}\exp(U).$ (2)
Our work examines the tradeoff between using the stronger property in (2) or increased capacity and the weaker property (1).
Figure 3: Pseudocode for the Lie Rotary Embedding (LieRE) attention mechanism. Unlike RoPE, LieRE allows for dense rotation matrices.
3.2 Attention Mechanism
LieRE is a modification of the standard attention mechanism to introduce positional information, which we review below. The modification we propose is independent of whether we use multiple heads, so we focus on single-headed attention for simplicity.
Let $X\in\mathbb{R}^{n\times d}$ be the set of input embeddings and $W_{Q},W_{K},W_{v}$ be learnable matrices. Let $Q=XW_{Q},K=XW_{K},V=XW_{V}$ be the keys, queries and values respectively. The outputs are computed as $\text{scores}=\frac{QK^{\top}}{\sqrt{d_{k}}}$, $\mathcal{W}=\text{softmax}(\text{scores})$ and final outputs $z=\mathcal{W}V$. We let $Q_{i}$ and $K_{i}$ denote the $i$th rows of $Q$ and $K$ respectively.
4 Method
LieRE is a simple modification to the attention mechanism that is presented in algorithm 3. Recall that we assume that positions are $n$-dimensional vectors, a matrix $A$ is skew-symmetric if $A^{T}=-A$, and that the matrix exponential of a skew-symmetric matrix, call it $A$, is always a high dimensional rotation matrix.
When encoding positions $p\in\mathbb{R}^{n}$, LieRE learns a skew-symmetric basis of matrices $\{A_{i}\}$ for $i\in[n]$. It encodes a position by writing it in this basis, $\sum\limits_{i=0}^{n}p_{i}A_{i}$. We then map the resulting skew-symmetric matrix to a high-dimensional rotation via the matrix exponential. $R(p)=\exp\left(\sum\limits_{i=0}^{n}p_{i}A_{i}\right)$ (Figure 1). Learning in the space of skew-symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices (Figure 2).
Figure 4: Pseudocode for the Rotary Position Embedding (RoPE) attention mechanism.
LieRE uses the rotation matrix computed above to modify the keys and queries of the standard attention mechanism. LieRE’s final step is to modify token $i$’s query and keys as $Q^{\prime}_{i}=R(p_{i})Q_{i}$ and $K^{\prime}_{i}=R(p_{i})K_{i}$. This modifies the score between tokens $i,j$ to be $X_{i}^{T}W_{Q}^{T}R(p_{i})^{T}R(p_{j})W_{K}X_{j}$. Recalling that $R^{T}=R^{-1}$ for any orthogonal matrix $R$ helps illustrate the encoding of relative positions in (1). Note that the only difference between LieRE and RoPE-Mixed is that the latter constrains the rotations to be block-diagonal with block size two.
We include the pseudocode for the LieRE attention in algorithm 3 in addition to the standard RoPE attention (algorithm 4). In practice, we compute the rotation matrices at the start of the forward pass. By default, the skew symmetric bases are learned separately for every layer and attention head except in the experimental section focused on sharing parameters across heads and layers.
Adjusting the skew-symmetric basis matrices’ block width allows us to incrementally adjust the capacity allocated towards position encoding. We specify the basis block width as a subscript, eg. LieRE8. When not specified, the block size is equal to the head dimension. If we set the block size to 2, we recover RoPE-Mixed (Heo et al., 2024). For a ViT-B model the head dimension is 64.
5 Experiments
In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022). We implement relative position encoding types including LieRE, Rope-Mixed, Visionllama, and absolute position encodings. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020). All experiments use RandAugment (Cubuk et al., 2020). We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B). We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively. Notably, a tile size of 2 corresponds to RoPE-Mixed.
5.1 2D Classification
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009). We further analyse different backbone models sizes and data ablation effects. We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024). For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (table 1).
Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results. All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
Method CIFAR-100 ImageNet-1k UCF101
Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 40.9 (40.5-41.3)
VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 45.0 (44.6-45.4)
RoPE-Mixed 68.8 (67.9-69.7) 68.8 (68.4-69.2) 46.3 (45.9-46.7)
LieRE8 70.3 (69.4-71.2) 69.6 (69.2-70.0) 47.0 (46.6-47.4)
LieRE64 70.0 (69.1-70.9) 69.3 (68.9-69.7) 44.7 (44.3-45.1)
We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020). Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in table 5. Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
Figure 5: Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions. Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 6).
Figure 6: Performance comparison of different positional embedding methods across varying dataset sizes. The plot shows the peak accuracy achieved by LieRE8, LieRE64, RoPE-Mixed, VisionLlaMA, and Absolute PE when trained on different fractions of the CIFAR-100 dataset. Both LieRE variants and RoPE-Mixed consistently outperform other methods, with their advantage becoming particularly pronounced in data-scarce scenarios.
5.2 Synthetic Spatial Classification
Figure 7: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down.
Recent observations indicate that even advanced models like ChatGPT-4 and Claude Sonnet 3.5 struggle with basic spatial reasoning tasks. To investigate whether these limitations stem from positional encoding mechanisms, we designed a synthetic image classification task (Shah et al., 2024). The task presents a $108\times 108$ pixel image containing a $9\times 9$ grid (81 cells). Within this grid, we randomly place eight arrows and six letters (A, B, C, D, E, and Y), ensuring that one arrow is placed in the direction of the base of the ”Y”. The objective is to identify the direction of this specific arrow. To introduce visual complexity, we include four spurious letters and seven additional arrows as distractors. Figure 7 illustrates an example of this task setup.
We verified that none of ChatGPT 4o, Claude Sonnet 3.5 and Gemeni Pro 1.5 are able to solve this task.
We train the models on 800,000 examples and observe that they generally converge after the first 400,000. The only exception to this is absolute position encodings, where we have variants trained on 800,000 and
Table 2: Comparing different positional encodings on synthetic task on base model (85M) equivalent to DeiT architecture across different resolutions, $108\times 108$, $168\times 168$, $276\times 276$
Method 108 168 276
Abs. Pos. E. 45.1 (43.1-47.7) 41.0 (39.0-43.2) 40.1 (38.2-42.0)
Abs. Pos. E. (2M examples) 47.2 (45.2-49.1) - -
VisionLlama RoPE 46.4 (44.0-47.9) 48.6 (46.0-49.9) 48.6 (46.2-50.1)
RoPE-Mixed 100 (99.5-100) 98.6 (97.6-98.7) 88.6 (87.4-89.9)
LieRE8 99.5 (99.2-99.8) 99.7 (99.4-99.9) 99.7 (99.5-99.9)
LieRE64 100 (99.5-100) 100 (99.6-100) 100 (99.7-100)
Table 2 outlines the performance of different positional encoding methods on a synthetic task using a base model (85M parameters) equivalent to the DeiT architecture. We evaluate the models across three different input resolutions ($108\times 108$, $168\times 168$, and $276\times 276$ pixels), revealing differences in scalability and effectiveness. While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently, suggesting superior resolution invariance of our Lie group-based approach. Qualitative analysis shows that absolute and visionllama position encoding attend less clearly to the ”Y” token than Rope-Mixed and LieRE. Please refer to the appendix for attention map examples, Figure 15 and Figure 15.
5.3 3D Classification
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024). For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012). Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022). The full set of hyperparameters may be found in appendix B. We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
5.4 LieRE parameter
LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is. To answer this question, we vary the density of the learned basis and the learnability across heads and layers.
We vary capacity by enforcing a block diagonal structure on the skew-symmetric basis. Varying the block size allows us to approach LieRE64. Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed. As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024). In figure 8 we evaluate accuracy versus block dimension (Appendix Table 8).
Figure 8: Performance varies with skew-symmetric basis learnable dimensions, referred to as LieRE parameters. This is equivalent to increasing the tile size in the skew-symmetric basis ($2\times 2,4\times 4,8\times 8,16\times 16,32\times 32,48\times 48$). For both 2D (CIFAR-100) and 3D (UCF101) LieRE with tile size $8\times 8$ performs superior.
Table 3: Accuracy with parameter sharing over heads and layers for ViT-B sized models on CIFAR-100.
FLOP All Shared Shared Across Layers Shared Across Heads RoPE-Mixed LieRE64 LieRE8
5.684G 68.8 70.0 70.3
5.684G 68.7 69.5 69.8
5.613G 69.5 69.7 69.7
5.613G 68.3 69.4 69.5
Table 3 presents the impact of sharing LieRE parameters across attention heads and layers on CIFAR-100 classification performance for RoPE-Mixed, LieRE64 and LieRE8 . We evaluate whether learning separate positional encodings for each attention head and layer provides performance benefits. We observe that learning across heads and layers yields superior performance followed by learning across layers.
5.5 Patch shuffling
Shuffling patches and frames allows us to see how much the model is able to use the positional information in its inputs. A model whose architecture does not allow/encourage the use of positional information will converge to a representation similar in spirit to a bag-of-words, where the relative locations of pixels/voxels do not matter. A greater dropoff in accuracy during shuffling is indicative that the model more heavily utilizes positional information.
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are more capable at using positional information. The complete results are displayed for CIFAR-100 and table for UCF101 (table 4).
Table 4: Relative accuracy drop for 2D image classification (CIFAR-100) and Video recognition (UCF101) after patch shuffling
CIFAR-100 (2D) UCF101 (3D)
Method Before After Drop(%) Before After Drop(%)
[…]
Abs. Pos. E. 63.9 19.6 69.3 40.9 39.5 0.0
VisionLlama RoPE 65.5 29.7 54.8 45.0 37.0 17.7
RoPE-Mixed 68.8 17.1 75.1 46.3 28.2 39.1
LieRE8 70.3 12.3 82.5 47.0 27.8 40.9
LieRE64 70.0 10.8 84.6 44.7 28.0 37.4
Figure 9: Validation accuracy comparison between LieRE 64 and RoPE-Mixed positional embeddings across different image resolutions on ImageNet. Left: Models trained at $224\times 224$ resolution. Right: Models pre-trained at $224\times 224$ and fine-tuned at $256\times 256$ resolution. Both approaches show similar performance trends up to $320\times 320$, after which LieRE 64 demonstrates significantly better accuracy retention at higher resolutions, particularly in the fine-tuned scenario.
5.6 Multi-resolution Classification
In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024). The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs. The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs. The full details can be found in appendix B. We evaluate the accuracy on the ImageNet validation set with varying inference resolutions. Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 9. For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions. For example, doubling the length of an image in each dimension doubles the range of positional indices. This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
6 Conclusion
In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities. Through experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101), we demonstrated that LieRE achieves better performance compared to existing positional encoding methods. Beyond improving accuracy, LieRE also exhibits data and compute efficiency. On CIFAR-100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings. Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency. The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities. By requiring no changes to the tokenizer other than outputting positions and no other code changes, LieRE may provide a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
7 Limitations
While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting. Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention. However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms. Future work could explore adaptations of our method to a wider range of architectures. Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$. While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics. Further research may be necessary to adapt our method to effectively handle such representational requirements.
Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
8 Broader Impacts
LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks. We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs. This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
Appendix A Appendix
Appendix B Experiment Hyperparameters
The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use a dropout of 0.1. We used CLS pooling in our implementation to facilitate comparability with existing literature in the field. Further experiments revealed substantial performance improvement with mean pooling and LieRE. We use the pytorch lightning framework for all experiments (Falcon, 2019).
B.1 2D Image Classification
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The basis capacity scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment (Cubuk et al., 2020). We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-Mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
B.2 3D Video Classifications
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
B.3 Multi-resolution Classification
The second training recipe consists of 30 epochs with an initial learning rate of 1E-5 with a cosine decay. This mirrors the DEIT III training reciple that first pretrains at a lower resolution and finetunes at a higher resolution.
B.4 CIFAR-100 Performance Across Model Scales
We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in table 5. We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
B.4.1 Basis parameters
Table 5: Comparison of Position Encoding Methods for Different ViT Models Sizes on CIFAR-100, Accuracy (bootstrapped 95%CI)
Position Enc. ViT-Tiny (22M) ViT-Base (86M) ViT-Large (302M)
Abs. Pos. E. 57.2 (56.2-58.1) 63.9 (62.9-65.8) 60.5 (59.5-61.4)
VisionLlaMA RoPE 58.2 (57.2-59.2) 65.5 (64.6 -66.5) 62.3 (60.4-64.2)
RoPE-Mixed 65.4 (64.5-66.4) 68.8 (67.9-69.7) 68.8 (67.9-69.7)
LieRE8 65.6 (64.7-66.6) 70.3 (69.4-71.2) 69.9 (68.9-70.8)
LieRE64 65.3 (64.4-66.3) 70.0 (69.1-69.7) 68.9 (68.0-69.8)
B.5 Python implementation of LieRE rotation matrix computation
1basis_raw_params = nn.Parameter(
2 torch.rand(
3 input_dimensionality,
4 head_dim,
5 head_dim,
6 ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper
7)
8
9upper_triangle = (
10 torch.triu(basis_raw_params, diagonal=1)
11)
12skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2)
13in_basis_positions = (
14 positions.reshape(list(positions.shape) + [1] * 2) * skew_bases
15)
16rotation_log = torch.sum(in_basis_positions, dim=-3)
17rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)).to(dtype=positions.dtype)
B.6 Compute efficiency
We demonstrate that LieRE-based transformer achieves comparable performance to the Absolute Position Embedding baseline (DeiT III (Touvron et al., 2022)) on CIFAR-100 with fewer training epochs. This represents a notable advancement over recent methods such as VisionLlama and RoPE-Mixed (Chu et al., 2024; Heo et al., 2024). Figure 10 illustrates that LieRE enables a 3.9$\times$ reduction in training compute while maintaining the accuracy achieved by absolute position encodings after 200 epochs.
Figure 10: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
B.7 FLOPS Comparison of methods
We find that since all methods we examine introduce a computational cost that is at most linear in the number of tokens, and runtime is dominated by the quadratic attention component, there is no substantial difference in computational efficiency between the methods. We list inference FLOP of the various methods in table 6.
Table 6: FLOP analysis with percentage increase compared to absolute position encodings
Position Enc. ViT-Tiny (22M) ViT-Base (85M) ViT-Large (302M)
Abs. Pos. E. 0.963G 5.607G 19.856G
VisionLlaMA RoPE 0.963G (+0.001%) 5.607G (+0.002%) 19.856G (+0.000%)
RoPE-Mixed 0.964G (+0.104%) 5.609G (+0.036%) 19.863G (+0.035%)
LieRE8 0.968G (+0.519%) 5.617G (+0.178%) 19.882G (+0.065%)
LieRE64 0.970G (+0.727%) 5.684G (+1.375%) 20.061G (+1.033%)
B.8 Validation Losses
Table 7: 2D image and 3D video classification Top-1 Validation loss (95% confidence intervals) results. All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
Method CIFAR-100 ImageNet-1k UCF101
Abs. Pos. E.∗,∗∗ 1.56 (1.47-1.56) 1.84 (1.81-1.86) 2.94 (2.92-2.96)
VisionLlama RoPE 1.56 (1.51-1.61) 1.98 (1.94-2.01) 2.66 (2.63-2.69)
RoPE-Mixed 1.38 (1.33-1.43) 1.72 (1.68-1.74) 2.52 (2.49-2.54)
LieRE8 1.36 (1.31-1.41) 1.73 (1.70-1.76) 2.47 (2.44-2.49)
LieRE64 1.37 (1.33-1.42) 1.73 (1.70-1.76) 2.64 (2.62-2.67)
B.9 Basis parameters
Table 8: Accuracy Results for Different $\mathrm{LieRE}_{\Theta}$ Parameters, relative to the model size of 85.2M and 88.7M
Dataset $\mathrm{LieRE}_{\Theta}$ Parameter Absolute $\mathrm{LieRE}_{\Theta}$ Parameter Relative Tile Size Accuracy (%) CI (95%)
CIFAR100 9216 0.01 % 2 68.84 (67.93-69.75)
CIFAR100 27648 0.03 % 4 69.28 (68.38-70.18)
CIFAR100 64512 0.08 % 8 70.32 (69.42-71.22)
CIFAR100 138240 0.16 % 16 69.85 (68.95-70.75)
CIFAR100 285696 0.34 % 32 69.65 (68.75-70.55)
CIFAR100 580608 0.68 % 64 69.99 (69.09-70.89)
UCF101 9216 0.01 % 2 46.30 (45.89-46.71)
UCF101 27648 0.03 % 4 45.67 (45.26-46.08)
UCF101 64512 0.08 % 8 47.03 (46.62-47.44)
UCF101 138240 0.16 % 16 46.86 (46.45-47.27)
UCF101 285696 0.32 % 32 46.42 (46.01-46.83)
UCF101 580608 0.66 % 64 44.68 (44.27-45.09)
B.10 Validation Accuracy Curves
Figure 11: CIFAR-100
B.11 Training Loss Curves
Figure 12: CIFAR-100
B.12 Attention Maps
We show the attention scores to the CLS token, averaged across heads for every layer. Each grid is rescaled so that the minimum element has value zero and maximum value of one. Red indicates the maximal score, and blue indicates the minimal value.
Figure 13: Attention Maps (normalized), $108\times 108$: Rope-Mixed, LieRE, Absolute and Visionllama positional encoding (x-axis), Layers 1-12 (y-axis). While Rope-Mixed and LieRE learn to look at the ”Y”, Absolute and Visionllama less and concentrate on the arrows.
Figure 14: Attention Maps (normalized), $276\times 276$: Rope-Mixed, LieRE, Absolute and Visionllama positional encoding (x-axis), Layers 1-12 (y-axis). While LieRE still learns to look at the ”Y”, Absolute and Visionllama do less so.
Figure 15: Comparison of attention maps across different resolutions and position encoding methods.